OpenStudy (anonymous):

Two equations are given below: m + 5n = 20 m = n − 4 What is the solution to the set of equations in the form (m, n)? (3, 7) (0, 4) (5, 1) (2, 6)

Vocaloid (vocaloid):

rewrite the first equation as m = 20 - 5n now we have m = 20 - 5n m = n - 4 we can set these two equations equal to each other 20 - 5n = n - 4 solve for n

OpenStudy (hybrik):

-6n=-24

OpenStudy (hybrik):

n = { # }

Vocaloid (vocaloid):

-6n = -24 divide both sides by (-6), what do you get?